OpenStudy (anonymous):

100 gm of KMnO3 are heated to redness. what volume of ocygen at 39 C and 765 atm pressure will evolve ? AnyOne .. explain please..

OpenStudy (mos1635):

KMnO3 or KMnO4 ????

OpenStudy (mos1635):

because if KMnO4 then 2 KMnO4 --Δ--> K2MnO4 + MnO2 + O2

OpenStudy (preetha):

Has to be KMnO4. There is no KMnO3

OpenStudy (preetha):

You have to set up the stoichiometry equation. Convert g of KMnO4 to moles of KMnO4. Then check the equation. And calculate the amount of Oxygen released with the 100 g (in moles) of KMnO4. Remember that 2 moles of KMnO4 gives you 1 mole of O2 gas. Then use PV=nRT and calculate the volume of oxygen evolved.

OpenStudy (anonymous):

>>>> sorry its KMnO4 ;)

OpenStudy (mos1635):

om go for it

OpenStudy (anonymous):

Can show me how to solve that please ?

OpenStudy (mos1635):

moles of KMnO4 = mass/MM n=100/(39+55+4*16) n=100/158 2 KMnO4 --Δ--> K2MnO4 + MnO2 + O2 2 moles produses 1 mole 100/158 mole produses 50/158 mole assume O2 ideal gas P*V=n*R*T V=n*R*T/P n=50/158 R=0.082 T=(39+273) P=765

OpenStudy (anonymous):

thanks (: